The data in the table to the right represent the calories and sugar​ (in grams) in one serving of seven different types of breakfast cereals.

a. Compute and interpret the coefficient of​ correlation, r.

b. At the 0.05 level of​ significance, is there a significant linear relationship between calories and​sugar?

Product   Calories   Sugar
Cereal 1   77   7
Cereal 2   102   2
Cereal 3   103   4
Cereal 4   110   4
Cereal 5   125   4
Cereal 6   186   10
Cereal 7   198   11
Answer #1

X: Calories

Y: Sugar

a) We have to find the coefficient of correlation r.

Formula is

x y xy x^2 y^2
77 7 539 5929 49
102 2 204 10404 4
103 4 412 10609 16
110 4 440 12100 16
125 4 500 15625 16
186 10 1860 34596 100
198 11 2178 39204 121
Total 901 42 6133 128467 322

n = 7

b)

Claim: There is a linear relationship between calories and sugar.

The null and alternative hypothesis is

Level of significance = 0.05

r = 0.7773

n = 7

Degrees of freedom = n - 2 = 7 - 2 = 5

Critical value = 0.754

( From pearson table)

r > critical value we reject null hypothesis.

Conclusion:

There is a linear relationship between calories and sugar.